# Architecture-Aware Learning Curve Extrapolation via Graph Ordinary Differential Equation

### v1

@@ IV-B Ablation Study
We explore three variations for each of the following components: the message passing mechanism, graph pooling, and time series encoder. The message passing mechanisms include GCN Kipf and Welling (2016), Graph Attention Networks (GAT) Veličković et al. (2017), and Graph Transformers Hu et al. (2020). For pooling methods, we use average pooling, max-pooling, and a learnable pooling method based on DiffPool Ying et al. (2018), where a separate GCN module determines each node’s contribution to the global embedding. The time-series encoder variations include self-attention Vaswani et al. (2017), TCN Pandey and Wang (2019), and an autoregressive version of GRU Dey and Salem (2017); Kipf et al. (2018). The model is trained using early stopping if no improvement is observed after 50 epochs.
@@ Appendix A Experimental Details
To generate MLP learning curves, we selected 2 tabular data classification tasks from OpenML Vanschoren et al. (2014) as detailed in Table IV. We adhere to the standard procedure outlined in LCBench Zimmer et al. (2021), with the exception that we introduce variability by randomly sampling the number of hidden units for each layer, rather than sampling only the maximal number of hidden units. All optimization runs are performed using SGD with decaying learning rate by 0.5 every 80 epochs. The total number of epochs is 200. We discard runs The parameters randomly sampled include four floating-point values and $2+l$ integers, where $l$ denotes the number of layers.
@@
Optimizer AdamW Loshchilov and Hutter (2017)
@@
The LSTM implementation is adapted from the published code associated with NRI Kipf et al. (2018). This implementation employs teacher forcing by utilizing the observation window. Specifically, it features a step() module, which comprises a Long-short-term memory (LSTM) block and a two-layer MLP with ReLU activation. The step() function processes the previous input state and hidden state, outputting the prediction and hidden state for the next immediate time step. The forward() function iteratively calls step() for a total of $T_{max}$ times. During the initial $T_{cond}$ steps, the observed learning curve data is used as the input state. For the subsequent steps beyond $T_{cond}$, the output from the previous prediction is used as the new input state.

### v3

@@ IV-B Ablation Study
We explore three variations for each of the following components: the message passing mechanism, graph pooling, and time series encoder. The message passing mechanisms include GCN (Kipf and Welling, 2016), Graph Attention Networks (GAT) (Veličković et al., 2017), and Graph Transformers (Hu et al., 2020). For pooling methods, we use average pooling, max-pooling, and a learnable pooling method based on DiffPool (Ying et al., 2018), where a separate GCN module determines each node’s contribution to the global embedding. The time-series encoder variations include self-attention (Vaswani et al., 2017), TCN (Pandey and Wang, 2019), and an autoregressive version of GRU (Dey and Salem, 2017; Kipf et al., 2018). The model is trained using early stopping if no improvement is observed after 50 epochs.
@@ Appendix A Experimental Details
To generate MLP learning curves, we selected 2 tabular data classification tasks from OpenML (Vanschoren et al., 2014) as detailed in Table IV. We adhere to the standard procedure outlined in LCBench (Zimmer et al., 2021), with the exception that we introduce variability by randomly sampling the number of hidden units for each layer, rather than sampling only the maximal number of hidden units. All optimization runs are performed using SGD with decaying learning rate by 0.5 every 80 epochs. The total number of epochs is 200. We discard runs The parameters randomly sampled include four floating-point values and $2+l$ integers, where $l$ denotes the number of layers.
@@
Optimizer AdamW (Loshchilov and Hutter, 2017)
@@
The LSTM implementation is adapted from the published code associated with NRI (Kipf et al., 2018). This implementation employs teacher forcing by utilizing the observation window. Specifically, it features a step() module, which comprises a Long-short-term memory (LSTM) block and a two-layer MLP with ReLU activation. The step() function processes the previous input state and hidden state, outputting the prediction and hidden state for the next immediate time step. The forward() function iteratively calls step() for a total of $T_{max}$ times. During the initial $T_{cond}$ steps, the observed learning curve data is used as the input state. For the subsequent steps beyond $T_{cond}$, the output from the previous prediction is used as the new input state.